OpenStudy (anonymous):

use the method of successive differences to determine the next term in the sequence: 4, 14, 82, 268, 657, 1359

OpenStudy (anonymous):

Example: 1,4,9,16,... find the first differences 4-1,9-4,16-9,...=3,5,7 find the second differences 5-3,7-5,...=2,2 keep doing this until u get until the level of difference where the terms are the same and not 0. 1 , 4 , 9 , 16 , x 3 , 5 , 7 , 9 2 , 2 , 2 x=16+9=25 so the next term is 25 for this example, try it with ur question

OpenStudy (anonymous):

also, the 9 in the first differences came from 7 + 2

OpenStudy (anonymous):

I am still confused

OpenStudy (anonymous):